# Multifidelity Simulation-based Inference for Computationally Expensive Simulators

### v3

@@ H.2 Low-fidelity model
H.3 Discussion on alternative solutions
@@
H.3.1 Pretraining on low fidelity samples
@@
H.3.2 Pretraining on the joint of LF and HF samples
@@
Appendix I Prior bounds across neuroscience tasks
@@
Appendix J Distance between the LF and HF posterior
@@
Appendix K Simulation versus training cost
We tracked the wall-clock run-time for training and simulation stages of the neural density estimator. Computations were performed on nodes each equipped with 4× Intel Xeon Gold 6448H CPUs (32 cores per socket, 128 physical cores, 256 logical CPUs) and approximately 2TB RAM, running Linux 5.14.0. We compare the costs in regimes where the performance of NPE is similar to MF-NPE and A-MF-TSNPE (Fig. 3). Details about the network architecture and hyperparameters are described in Appendix D.1. In cases where many samples had to be generated for active non-amortized schemes (e.g., $10^{5}$ HF samples for the L5PC task; Figure 3), we used multiprocessing over CPUs. The simulations for the third task were parallelized over 913 CPUs.
@@
Input: Simulations $\{(\boldsymbol{\theta},\boldsymbol{x}^{(f)})\}^{F}_{f=1}$ over $F$ fidelities; Early stopping criterion $S$; conditional density estimators $\{q_{\psi}^{(f)}(\boldsymbol{\theta}|\boldsymbol{x}^{(f)})\}^{F}_{f=1}$ with features $\psi$.
for $f=1$ to $F$ do
  $\mathcal{L}(\psi^{(f)})=\frac{1}{N^{(f)}}\sum_{i=1}^{N^{(f)}}-\log q^{(f)}_{\psi}\left(\boldsymbol{\theta}_{i}|\boldsymbol{x}^{(f)}_{i}\right)$ .
  $\mathrm{opt}^{(f)}\leftarrow\mathrm{Adam}(\cdot)$
  if $f>1$ then
   Initialize $q_{\psi}^{(f)}$ with features of trained $q_{\psi}^{(f-1)}$.
  end if
  for epoch in epochs do
   train $q_{\psi}^{(f)}$ to minimize $\mathcal{L}(\psi^{(f)})$ until $S$ is reached.
  end for
end for
@@
Input: NN pairs of $(\boldsymbol{\theta},\boldsymbol{x}_{\text{L}})$; conditional density estimators $q_{\psi}(\boldsymbol{\theta}|\boldsymbol{x}_{\text{L}})$ and $q_{\phi}(\boldsymbol{\theta}|\boldsymbol{x})$ with learnable parameters $\psi$ and $\phi$; early stopping criterion $S$; simulator $p_{\mathrm{H}}(\boldsymbol{x}|\boldsymbol{\theta})$; prior $p(\boldsymbol{\theta})$; number of rounds $R$; $\epsilon$ that defines the highest-probability region ($\mathrm{HPR}_{\epsilon}$); number of high-fidelity simulations per round $M$.
Output: posterior estimate $q_{\phi}(\boldsymbol{\theta}|\boldsymbol{x})$
$\mathcal{L}(\psi)=\frac{1}{N}\sum_{i=1}^{N}-\log q_{\psi}\left(\boldsymbol{\theta}_{i}|\boldsymbol{x}^{\text{L}}_{i}\right)$ .
for epoch in epochs do
  train $q_{\psi}$ to minimize $\mathcal{L(\psi)}$ until $S$ is reached.
end for
 Initialize $\tilde{p}(\boldsymbol{\theta})$ as $p(\boldsymbol{\theta})$
 Initialize $q_{\phi}$ with weights and biases of trained $q_{\psi}$.
for r in $R$ do
  $\boldsymbol{\theta}^{(r)}\sim\tilde{p}(\boldsymbol{\theta})$, sample parameters from proposal
  $\boldsymbol{x}^{(r)}\sim p_{\mathrm{H}}(\boldsymbol{x}|\boldsymbol{\theta}^{(r)})$, generate high-fidelity simulations
  for epoch in epochs do
   $\mathcal{L}(\phi)=\frac{1}{M}\sum_{i=1}^{M}-\log q_{\phi}\left(\boldsymbol{\theta}^{(r)}_{i}|\boldsymbol{x}^{(r)}_{i}\right)$.
   train $q_{\phi}$ to minimize $\mathcal{L(\phi)}$ until $S$ is reached.
  end for
  Compute expected coverage ($\tilde{p}(\boldsymbol{\theta})$, $q_{\phi}$)
@@
  $\tilde{p}(\boldsymbol{\theta})\propto p(\boldsymbol{\theta})\cdot\mathbbm{1}_{\theta\in\mathrm{HPR}_{\epsilon}}$
end for
@@
All experiments were ran with $R=5$ rounds and $\epsilon=1e^{-6}$. More details about TSNPE at Deistler et al. (2022).
M.2 A-MF-TSNPE
Algorithm 4 A-MF-TSNPE
@@
Input: NN pairs of $(\boldsymbol{\theta},\boldsymbol{x}_{\text{L}})$; conditional density estimator $q_{\psi}(\boldsymbol{\theta}|\boldsymbol{x}_{\text{L}})$ with learnable parameters $\psi$ and and ensemble of conditional density estimators $\{q^{e}_{\phi}(\boldsymbol{\theta}|\boldsymbol{x})\}^{e}_{E}$, each with independent $\phi$; early stopping criterion $S$; simulator $p_{\mathrm{H}}(\boldsymbol{x}|\boldsymbol{\theta})$; prior $p(\boldsymbol{\theta})$; number of rounds $R$; $\epsilon$ that defines the highest-probability region ($\mathrm{HPR}_{\epsilon}$); number of high-fidelity simulations per round $M$.
Output: Ensemble posterior estimate $q_{\phi}(\boldsymbol{\theta}|\boldsymbol{x})=\frac{1}{E}\sum^{E}_{e=1}q^{e}_{\phi}(\boldsymbol{\theta}|\boldsymbol{x})$
$\mathcal{L}(\psi)=\frac{1}{N}\sum_{i=1}^{N}-\log q_{\psi}\left(\boldsymbol{\theta}_{i}|\boldsymbol{x}^{\text{L}}_{i}\right)$ .
for epoch in epochs do
  train $q_{\psi}$ to minimize $\mathcal{L(\psi)}$ until $S$ is reached.
end for
for e $\in$ Ensemble do
  Initialize $q_{\phi}^{e}$ with weights and biases of trained $q_{\psi}$.
end for
$\boldsymbol{\theta}_{\mathrm{pool}}\sim p(\boldsymbol{\theta})$
 Initialize $\tilde{p}(\boldsymbol{\theta})$ as $p(\boldsymbol{\theta})$
for r in $R$ do
@@
  $\boldsymbol{\theta}^{(r)}_{\mathrm{prop}}\sim\tilde{p}(\boldsymbol{\theta})$, generate MBM-B samples from proposal
@@
  $\boldsymbol{\theta}^{(r)}_{\mathrm{active}}$ = top $B$ values from $\boldsymbol{\theta}_{\mathrm{pool}}$ using the acquisition function eq. (2)
  $\boldsymbol{\theta}^{(r)}=\{\boldsymbol{\theta}_{\mathrm{prop}}^{(r)}\cup\boldsymbol{\theta}_{\mathrm{active}}^{(r)}\}$
  $\boldsymbol{x}^{(r)}\sim p_{\mathrm{H}}(\boldsymbol{x}|\boldsymbol{\theta}^{(r)})$, generate high-fidelity simulations
  for e $\in$ Ensemble do
   for epoch in epochs do
    $\mathcal{L}(\phi)=\frac{1}{M}\sum_{i=1}^{M}-\log q^{e}_{\phi}\left(\boldsymbol{\theta}^{(r)}_{i}|\boldsymbol{x}^{(r)}_{i}\right)$.
    train $q_{\phi}$ to minimize $\mathcal{L(\phi)}$ until $S$ is reached.
   end for
  end for
  Compute expected coverage ($\tilde{p}(\boldsymbol{\theta})$, $\frac{1}{E}\sum q^{e}_{\phi}(\boldsymbol{\theta}|\boldsymbol{x})$)
@@
  $\tilde{p}(\boldsymbol{\theta})\propto p(\boldsymbol{\theta})\cdot\mathbbm{1}_{\theta\in\mathrm{HPR}_{\epsilon}}$
end for
@@
All experiments were ran with $R=5$ rounds, $\epsilon=1e^{-6}$, and an ensemble of 5 networks. The addition of an acquisition function biases the proposal distribution, causing the density estimate to diverge from the true posterior. In principle, this could be addressed by using atomic proposals (Greenberg et al., 2019), but given that such an approach suffers from posterior leakage, we do not introduce a proposal correction in order to retain the well-behaved loss function in TNSPE. We argue that the benefit of informative samples would outweigh the potential bias, as long as the percentage of samples selected from the acquisition function would be small compared to the proposal samples. Therefore, we set $B=.2M$ to mitigate the concern of biasing the posterior with parameters selected with the acquisition function.

### v4

@@ G.2 Low-fidelity model
G.3 Synaptic Plasticity with varying parameter space
We investigated how inference performance changes as the discrepancy between the low- and high-fidelity models increases. To this end, we varied the dimensionality of the low-fidelity model between 3, 6, and 12 parameters, while keeping the high-fidelity model fixed at 24 parameters. Parameters that were excluded from inference in the low-fidelity settings were fixed to the following values for each connection type: $\tau_{\text{pre}}=\tau_{\text{post}}=0.05,\gamma=-1.9,\alpha=\beta=\kappa=0.5$. The value of $\gamma$ should be smaller than other parameters to fulfill the second-order stability condition (Confavreux et al., 2023).
Figure 19: Negative-log-likelihood over true parameters, with different numbers of free parameters in the low-fidelity model.
We observe that the performance of MF-NPE degrades as the number of parameters in the low-fidelity model decreases as compared to the high-fidelity model. In particular, unlike in all our other experiments, when the low-fidelity model had only 3 parameters, pretraining on $10^{5}$ low-fidelity samples led to worse MF-NPE performance: in this regime, using $10^{5}$ samples (MF-NPE5) resulted in negative transfer, whereas pretraining on $10^{4}$ samples (MF-NPE4) resulted in a performance close to standard NPE.
G.4 Discussion on alternative solutions
@@
G.4.1 Pretraining on low fidelity samples
@@
G.4.2 Pretraining on the joint of LF and HF samples
@@
Appendix H Prior bounds across neuroscience tasks
@@
Appendix I Distance between the LF and HF posterior
@@ I.1 Pairplots
Appendix J Simulation versus training cost
We tracked the wall-clock run-time for training and simulation stages of the neural density estimator. Computations were performed on nodes each equipped with 4× Intel Xeon Gold 6448H CPUs (32 cores per socket, 128 physical cores, 256 logical CPUs) and approximately 2TB RAM, running Linux 5.14.0. We compare the costs in regimes where the performance of NPE is similar to MF-NPE and MF-TSNPE-AF (Fig. 3). Details about the network architecture and hyperparameters are described in Appendix C.1. In cases where many samples had to be generated for active non-amortized schemes (e.g., $10^{5}$ HF samples for the L5PC task; Figure 3), we used multiprocessing over CPUs. The simulations for the third task were parallelized over 913 CPUs.
@@
Appendix K TARP evaluation for all tasks
We performed additional evaluations on the calibration of all experiments with TARP (Lemos et al., 2023).
Figure 22: TARP calibration test across $10^{5}$ LF simulations ($10^{4}$ for the Gaussian blob example). The calibration test was performed over 200 runs.
@@
1:Input: Simulations $\{(\boldsymbol{\theta},\boldsymbol{x}^{(f)})\}^{F}_{f=1}$ over $F$ fidelities; Early stopping criterion $S$; conditional density estimators $\{q_{\psi}^{(f)}(\boldsymbol{\theta}|\boldsymbol{x}^{(f)})\}^{F}_{f=1}$ with features $\psi$.
2:for $f=1$ to $F$ do
3:  $\mathcal{L}(\psi^{(f)})=\frac{1}{N^{(f)}}\sum_{i=1}^{N^{(f)}}-\log q^{(f)}_{\psi}\left(\boldsymbol{\theta}_{i}|\boldsymbol{x}^{(f)}_{i}\right)$ .
4:  $\mathrm{opt}^{(f)}\leftarrow\mathrm{Adam}(\cdot)$
5:  if $f>1$ then
6:   Initialize $q_{\psi}^{(f)}$ with features of trained $q_{\psi}^{(f-1)}$.
7:  end if
8:  for epoch in epochs do
9:   train $q_{\psi}^{(f)}$ to minimize $\mathcal{L}(\psi^{(f)})$ until $S$ is reached.
10:  end for
11:end for
@@
1:Input: NN pairs of $(\boldsymbol{\theta},\boldsymbol{x}_{\text{L}})$; conditional density estimators $q_{\psi}(\boldsymbol{\theta}|\boldsymbol{x}_{\text{L}})$ and $q_{\phi}(\boldsymbol{\theta}|\boldsymbol{x})$ with learnable parameters $\psi$ and $\phi$; early stopping criterion $S$; simulator $p(\boldsymbol{x}|\boldsymbol{\theta})$; prior $p(\boldsymbol{\theta})$; number of rounds $R$; $\epsilon$ that defines the highest-probability region ($\mathrm{HPR}_{\epsilon}$); number of high-fidelity simulations per round $M$.
2:Output: posterior estimate $q_{\phi}(\boldsymbol{\theta}|\boldsymbol{x})$
[…]
4:for epoch in epochs do
5:  train $q_{\psi}$ to minimize $\mathcal{L(\psi)}$ until $S$ is reached.
6:end for
7: Initialize $\tilde{p}(\boldsymbol{\theta})$ as $p(\boldsymbol{\theta})$
8: Initialize $q_{\phi}$ with weights and biases of trained $q_{\psi}$.
9:for r in $R$ do
10:  $\boldsymbol{\theta}^{(r)}\sim\tilde{p}(\boldsymbol{\theta})$, sample parameters from proposal
11:  $\boldsymbol{x}^{(r)}\sim p(\boldsymbol{x}|\boldsymbol{\theta}^{(r)})$, generate high-fidelity simulations
12:  for epoch in epochs do
13:   $\mathcal{L}(\phi)=\frac{1}{M}\sum_{i=1}^{M}-\log q_{\phi}\left(\boldsymbol{\theta}^{(r)}_{i}|\boldsymbol{x}^{(r)}_{i}\right)$.
14:   train $q_{\phi}$ to minimize $\mathcal{L(\phi)}$ until $S$ is reached.
15:  end for
16:  Compute expected coverage ($\tilde{p}(\boldsymbol{\theta})$, $q_{\phi}$)
@@
17:  $\tilde{p}(\boldsymbol{\theta})\propto p(\boldsymbol{\theta})\cdot\mathbbm{1}_{\theta\in\mathrm{HPR}_{\epsilon}}$
18:end for
@@
All experiments were run with $R=5$ rounds and $\epsilon=1e^{-6}$. More details about TSNPE at Deistler et al. (2022).
M.2 MF-TSNPE-AF
Algorithm 4 MF-TSNPE-AF
@@
1:Input: NN pairs of $(\boldsymbol{\theta},\boldsymbol{x}_{\text{L}})$; conditional density estimator $q_{\psi}(\boldsymbol{\theta}|\boldsymbol{x}_{\text{L}})$ with learnable parameters $\psi$ and ensemble of conditional density estimators $\{q^{e}_{\phi}(\boldsymbol{\theta}|\boldsymbol{x})\}^{e}_{E}$, each with independent $\phi$; early stopping criterion $S$; simulator $p(\boldsymbol{x}|\boldsymbol{\theta})$; prior $p(\boldsymbol{\theta})$; number of rounds $R$; $\epsilon$ that defines the highest-probability region ($\mathrm{HPR}_{\epsilon}$); number of high-fidelity simulations per round $M$.
2:Output: Ensemble posterior estimate $q_{\phi}(\boldsymbol{\theta}|\boldsymbol{x})=\frac{1}{E}\sum^{E}_{e=1}q^{e}_{\phi}(\boldsymbol{\theta}|\boldsymbol{x})$
[…]
4:for epoch in epochs do
5:  train $q_{\psi}$ to minimize $\mathcal{L(\psi)}$ until $S$ is reached.
6:end for
7:for e $\in$ Ensemble do
8:  Initialize $q_{\phi}^{e}$ with weights and biases of trained $q_{\psi}$.
9:end for
[…]
11: Initialize $\tilde{p}(\boldsymbol{\theta})$ as $p(\boldsymbol{\theta})$
12:for r in $R$ do
@@
13:  $\boldsymbol{\theta}^{(r)}_{\mathrm{prop}}\sim\tilde{p}(\boldsymbol{\theta})$, generate MBM-B samples from proposal
@@
14:  $\boldsymbol{\theta}^{(r)}_{\mathrm{active}}$ = top $B$ values from $\boldsymbol{\theta}_{\mathrm{pool}}$ using the acquisition function (2)
15:  $\boldsymbol{\theta}^{(r)}=\{\boldsymbol{\theta}_{\mathrm{prop}}^{(r)}\cup\boldsymbol{\theta}_{\mathrm{active}}^{(r)}\}$
16:  $\boldsymbol{x}^{(r)}\sim p(\boldsymbol{x}|\boldsymbol{\theta}^{(r)})$, generate high-fidelity simulations
17:  for e $\in$ Ensemble do
18:   for epoch in epochs do
19:    $\mathcal{L}(\phi)=\frac{1}{M}\sum_{i=1}^{M}-\log q^{e}_{\phi}\left(\boldsymbol{\theta}^{(r)}_{i}|\boldsymbol{x}^{(r)}_{i}\right)$.
20:    train $q_{\phi}$ to minimize $\mathcal{L(\phi)}$ until $S$ is reached.
21:   end for
22:  end for
23:  Compute expected coverage ($\tilde{p}(\boldsymbol{\theta})$, $\frac{1}{E}\sum q^{e}_{\phi}(\boldsymbol{\theta}|\boldsymbol{x})$)
@@
24:  $\tilde{p}(\boldsymbol{\theta})\propto p(\boldsymbol{\theta})\cdot\mathbbm{1}_{\theta\in\mathrm{HPR}_{\epsilon}}$
25:end for
@@
All experiments were run with $R=5$ rounds, $\epsilon=1e^{-6}$, and an ensemble of 5 networks. The addition of an acquisition function biases the proposal distribution, causing the density estimate to diverge from the true posterior. In principle, this could be addressed by using atomic proposals (Greenberg et al., 2019), but given that such an approach suffers from posterior leakage, we do not introduce a proposal correction in order to retain the well-behaved loss function in TSNPE. We argue that the benefit of informative samples would outweigh the potential bias, as long as the percentage of samples selected from the acquisition function would be small compared to the proposal samples. Therefore, we set $B=.2M$ to mitigate the concern of biasing the posterior with parameters selected with the acquisition function.